OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –6 –6 –4 –3 Line 2 x y 0 3 2 6 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (anonymous):

i am 99.99% percent sure that it is 0!

OpenStudy (anonymous):

i am just not sure on the 1% part i need someone to check plz don't guess

OpenStudy (anonymous):

@livias.random whenever u r ready can u check?

OpenStudy (anonymous):

@livias.random

OpenStudy (anonymous):

i am writing down my answer to check if the 0 is right i hope it is!

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

oops i meant i am checking my equation to see if it is 0

OpenStudy (anonymous):

ok i got 0

OpenStudy (anonymous):

what did u get?

OpenStudy (anonymous):

m = (-6-(-3))/(-6-(-4)) m = -3/-2 m = 3/2 (y-(-6)) = 3/2(x-(-6)) y+6 = 3/2x + 18/2 y = 3/2x + 3 - m = (3-(6))/(0-2) m = -3/-2 m = 3/2 (y-3) = 3/2(x-0) y-3 = 3/2x y = 3/2x + 3 - 3/2x+3 = 3/2x+3 so x = x so any value of x will work so D. Infinitely many

OpenStudy (anonymous):

omg really? i am so re checking my equation one sec

OpenStudy (anonymous):

You like Animal Jam too?

OpenStudy (anonymous):

@Houdini_Dragon

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Awesome!